aliha:

What is the mass of a stone moving at a speed of 15 meters/second and having a momentum of 7.5 kilogram·meters/second? A. 0.5 kilograms B. 1.5 kilograms C. 11 kilograms D. 1.1 × 102 kilograms

Razor:

The formula for this is; p = m × v where, p = momentum m = mass of object v = velocity of object p = 7.5 x 15 To calculate the mass of the stone we go in this direction > \[m = \frac{ p }{ v} = \frac{ 7.5kgm/s }{ 15m/s} = 0.5kg\]

Razor:

So your answer would be 0.5kg